OpenStudy (anonymous):

T-Formules: How does one go from the system \[sin(2a)=2sin(a)*cos(a)\] and \[sin^2(a)+cos^2(a)=1\] to \[sin(2a)=\frac{2sin(a)*cos(a)}{sin^2(a)+cos^2(a)}\]

OpenStudy (anonymous):

\[\sin(2a)=\frac{2\sin(a)*\cos(a)}{\sin^2(a)+\cos^2(a)}\]

OpenStudy (cggurumanjunath):